# A Unified Architecture for N-Dimensional Visualization and Simulation: 4D Implementation and Evaluation including Boolean Operations

### v3

@@ Abstract
This paper proposes a unified software architecture for visualization and simulation based on a design targeting an $N$-dimensional space. The contribution of this work lies in presenting an architectural configuration that integrates multiple processes into a single software architecture: Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (including orientation and view transformations), and hyperplane slicing for visualization. The proposed approach adopts an approximate implementation that tolerates numerical errors and prioritizes implementation transparency over guarantees of numerical rigor. The experimental results and evaluations presented in this paper are limited to a 4D implementation; no evaluation is conducted for $N>4$, and the discussion is restricted to stating that the architecture itself has a dimension-independent structure. This paper also proposes an interaction design for high-dimensional exploration based on FPS navigation. As an input example involving shape changes over time, a non-rigid body simulation based on XPBD (Extended Position Based Dynamics) is integrated into the 4D implementation. Experimental results confirm that the 4D implementation runs on a single PC.
@@ 1 Introduction
This paper builds on prior work on visualization and simulation in high-dimensional spaces. However, examples remain limited that integrate Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (orientation and view transformations), and hyperplane slicing for visualization into a single software architecture, under the assumption of an iterative workflow of user actions and result inspection. In this work, we propose a unified architecture intended for use by researchers and developers. While this architecture is designed for an $N$-dimensional space, the present study implements and evaluates it in 4D.
The contribution of this work is to present an architectural configuration that integrates the above processing components as a single software architecture and to evaluate the 4D implementation experimentally.
In this paper, interactivity is defined as a workflow that presupposes repeated cycles of user actions and result inspection, in contrast to offline batch processing in which input conditions are fixed in advance and executed in a single run. Hereafter, the term “interactive,” including its adjectival and adverbial usages, refers to usage scenarios that satisfy or require this notion of interactivity.
@@ 3.6.2 Facet Clipping Algorithm in the 4D Implementation
To generate 3D cross-sections of objects embedded in 4D space, the present implementation computes the geometric intersections between each facet of the 4D mesh, represented as tetrahedra, and an axis-aligned hyperplane $w=w_{\text{slice}}$, where $w_{\text{slice}}$ is an arbitrary constant. This computation is performed for all facets of the input mesh. Cross-section generation based on this type of slicing plane has also been adopted in existing studies on interactive 4D visualization [5]. In this implementation, no mechanism is provided to arbitrarily specify the orientation of the cutting hyperplane; cross-sections are always generated using the axis-aligned plane $w=w_{\text{slice}}$. When observing the cross-sections from arbitrary directions, changes in the viewing direction are handled as view transformations.
@@ 3.8 Design Based on Topological and Geometric Separation
In the 4D implementation, based on this design, a non-rigid body simulation based on XPBD is integrated, and it is confirmed that mesh vertex coordinates are updated over time and reflected in cross-section generation and rendering. However, evaluation of the validity or accuracy of the physical model is not included in the scope of this study.
@@ 4.1 Background and Challenges
In the $N$-dimensional simulation environments targeted by this study, there are cases where a means for exchanging high-dimensional mesh data between projects is required. However, for $N$-dimensional meshes, there are only a limited number of common exchange formats that are widely used across the research community. In practice, project-specific formats or extensions of existing formats are often employed. This situation can lead to additional effort when sharing data across different implementations or research groups, including data conversion, validation, and interpretation of specifications.
To address this issue, this study defines a chunk-based file format, “Plex” (.plex), for the purpose of exchanging $N$-dimensional mesh data.
@@ 5.2.2 Boolean Operation Performance
The measurement results are shown in Table 4. In 3D space, for the combinations tested, the computation time was generally on the order of several tens of milliseconds. In contrast, in 4D space the computation time increased, reaching approximately $300\text{\,}\mathrm{ms}$ for Cube–Cube cases and exceeding $2\text{\,}\mathrm{s}$ for Random–Random cases in some instances.
@@ 5.5 Discussion and Limitations
6 Conclusion and Future Work
In this paper, we proposed a unified architecture for visualization and simulation based on a design targeting $N$-dimensional space. We presented a configuration that integrates multiple processes, including convex hull mesh generation, Boolean operations, coordinate transformations, and hyperplane slicing, into a single software architecture, and adopted the separation of topology and geometry as a design principle. In the 4D implementation, a non-rigid body simulation based on XPBD was implemented within the same architecture, and it was confirmed that updates to mesh vertex coordinates are reflected in cross-section generation and rendering. In addition, an interaction scheme for high-dimensional exploration was designed, and implemented in the 4D implementation.
Furthermore, as one example of data representation for $N$-dimensional meshes, this study designed the Plex data format and defined its specification.
@@ 6 Conclusion and Future Work
Future work includes examining the impact of increased computational cost due to the growth of combinatorial complexity in intersection tests and simplicial subdivision for algorithms in five or more dimensions; investigating more numerically robust and physically accurate $N$-dimensional non-rigid physical models; exploring integration using other phenomena, such as fluid simulations, as input examples; and considering GPU implementations for components that are suitable for parallelization among the processes implemented in a CPU-based manner in this paper. In addition, conducting comparative evaluations with existing computational geometry libraries and previously reported methods, and quantitatively characterizing the trade-offs among computation time, memory usage, and robustness to degenerate configurations, also remain as future tasks. Furthermore, conducting qualitative evaluations also remains as future work. Such evaluations include collecting structured feedback from users and domain researchers on representative exploration and editing tasks.

### v4

@@ Abstract
This paper proposes a unified software architecture for visualization and simulation based on a design targeting an $N$-dimensional space. The contributions of this study are twofold. First, it presents an architectural configuration that integrates multiple processes into a single software architecture: Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (pose transformation and view transformation), and hyperplane slicing for visualization. Second, it defines “Plex” (.plex) as a file format intended for the exchange of $N$-dimensional mesh data. The proposed approach adopts an approximate implementation that tolerates numerical errors and prioritizes implementation transparency over guarantees of numerical rigor. The experimental results and evaluations presented in this paper are limited to a 4D implementation; no evaluation is conducted for $N>4$, and the discussion is restricted to stating that the architecture itself has a dimension-independent structure. This paper also proposes an interaction design for high-dimensional exploration based on FPS navigation. As an input example involving shape changes over time, a non-rigid body simulation based on XPBD (Extended Position Based Dynamics) is integrated into the 4D implementation. Experimental results confirm that the 4D implementation runs on a single PC.
@@ 1 Introduction
This paper builds on prior work on visualization and simulation in high-dimensional spaces. However, examples remain limited that integrate Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (orientation and view transformations), and hyperplane slicing for visualization into a single software architecture, under the assumption of an iterative workflow of user actions and result inspection.
This study proposes a unified architecture intended for use by researchers and developers. When exchanging high-dimensional mesh data across projects, the availability of common formats is limited. Consequently, project-specific formats or extensions of existing formats are often employed.
This study has two main contributions. First, it presents a configuration that integrates the above processing components into a single software architecture and evaluates a four-dimensional implementation through experiments. Second, it defines “Plex” (.plex) as a file format intended for the exchange of $N$-dimensional mesh data.
The architecture is designed for $N$-dimensional space; however, the implementation and evaluation in this study are limited to 4D. In this paper, interactivity is defined as a workflow that presupposes repeated cycles of user actions and result inspection, in contrast to offline batch processing in which input conditions are fixed in advance and executed in a single run. Hereafter, the term “interactive,” including its adjectival and adverbial usages, refers to usage scenarios that satisfy or require this notion of interactivity.
@@ 3.6.2 Facet Clipping Algorithm in the 4D Implementation
To generate 3D cross-sections of objects embedded in 4D space, the present implementation computes the geometric intersections between each facet of the 4D mesh, represented as tetrahedra, and an axis-aligned hyperplane $w=w_{\text{slice}}$, where $w_{\text{slice}}$ is an arbitrary constant. This computation is performed for all facets of the input mesh. Cross-section generation based on this type of slicing plane has also been adopted in existing studies on interactive 4D visualization [5]. Since rotating the object’s orientation via view transformation enables cross-sectional observation from arbitrary directions, this implementation does not employ a procedure to arbitrarily specify the orientation of the cutting hyperplane. Cross-sections are always generated using the axis-aligned plane $w=w_{\text{slice}}$.
@@ 3.8 Design Based on Topological and Geometric Separation
In the 4D implementation, based on this design, a non-rigid body simulation based on XPBD is integrated, and it is confirmed that mesh vertex coordinates are updated over time and reflected in cross-section generation and rendering. This serves as an implementation example demonstrating that new functionality can be integrated in the four-dimensional implementation without modifying the existing logic. Here, the role of XPBD is limited to serving as an example of a data source that provides dynamically changing geometry, and the validity or accuracy of the physical model is not within the scope of this study.
@@ 4.1 Background and Challenges
As described above, when exchanging high-dimensional mesh data across projects, the availability of common formats is limited. This situation can lead to additional effort when sharing data across different implementations or research groups, including data conversion, validation, and interpretation of specifications.
Since the architecture proposed in this study is intended for use by researchers and developers, we define a chunk-based file format, “Plex” (.plex), to enable the exchange of $N$-dimensional mesh data.
@@ 5.2.2 Boolean Operation Performance
The measurement results are shown in Table 4. In 3D space, for the combinations tested, the computation time was generally on the order of several tens of milliseconds. In contrast, in 4D space the computation time increased, reaching approximately $300\text{\,}\mathrm{ms}$ for Cube–Cube cases and exceeding $2\text{\,}\mathrm{s}$ for Random–Random cases in some instances. Numerical-error-induced holes or cracks in the mesh were not observed by visual inspection.
@@ 5.5 Discussion and Limitations
6 Conclusion
In this paper, we proposed a unified architecture for visualization and simulation based on a design targeting $N$-dimensional space. We presented a configuration that integrates multiple processes, including convex hull mesh generation, Boolean operations, coordinate transformations, and hyperplane slicing, into a single software architecture, and adopted the separation of topology and geometry as a design principle. In addition, an interaction scheme for high-dimensional exploration was designed. Furthermore, as one example of data representation for $N$-dimensional meshes, this study designed the Plex data format and defined its specification.
In the 4D implementation, a non-rigid body simulation based on XPBD was implemented within the same architecture, and it was confirmed that updates to mesh vertex coordinates are reflected in cross-section generation and rendering.
7 Future Work
@@
Future work includes examining the impact of increased computational cost due to the growth of combinatorial complexity in intersection tests and simplicial subdivision for algorithms in five or more dimensions; ensuring the topological correctness of meshes generated by Direct Quickhull; investigating more numerically robust and physically accurate $N$-dimensional non-rigid physical models; exploring integration using other phenomena, such as fluid simulations, as input examples; and considering GPU implementations for components that are suitable for parallelization among the processes implemented in a CPU-based manner in this paper. In addition, conducting comparative evaluations with existing computational geometry libraries and previously reported methods, and quantitatively characterizing the trade-offs among computation time, memory usage, and robustness to degenerate configurations, also remain as future tasks. Furthermore, conducting qualitative evaluations also remains as future work. Such evaluations include collecting structured feedback from users and domain researchers on representative exploration and editing tasks.